APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Thursday February 2, 2006 (MJD 53769) ---=== OBSERVING TEAM ===--- Swing: K. Pan Night: V. Malanuskenko, M. Harvanek, E. Malanuskenko ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- Nearly all of the night was lost to bad weather and poor observing conditions. We obtained 3 usable science exposures and 6 bad ones. The night was spent fighting poor seeing, clouds and high wind. Useful data obtained for 2 plates (1245, 2387). Brian Yanny and Peregrine: If you're bored you might want to look at some of the data we declared bad on plates 2387 and 1245, respectively. Maybe you'll find something worth salvaging. ---=== OBSERVING LOG ===--- Afternoon: ---------- iccSoundOff returned ad1 was not available. Astro crates were rebooted. Spectrograph focus checked and no change made. the best focus temp is around 2 C. As usual, details in problem section. 22:50Z Fans on, louvers open. 23:20Z east door open and west door cracked. cart 8 (plt 1245) mounted. Night: ------ 00:45Z - Enclosure off. Partly cloudy so we begin with spectroscopy (plate 1245, cartridge 8). However, before we can even get set up on the field, the clouds thicken. Radar shows returns to the north coming toward us and we can see virga over the basin. So, we decide to close. 01:20Z - Enclosure on, doors partially closed, fans still on. The threatening clouds moved past and we opened again. 03:00Z - Enclosure off. Still some clouds so we stay with spectroscopy. 03:09Z - Plate 1245 (cartridge 8): Orion. Guide stars on the gotoField. A record of the exposures for this plate is given below. Note that the seeing ranged from poor to awful. This, combined with moon and clouds, made it difficult to obtain useful data and exposures 37003 and 37004 were both declared bad. Thus, we got only one good science/sky pair (900 sec each) before this plate expired. exposure # exposure time purpose guided seeing or offset ---------- ------------- ------- ------ ---------------- 37001 900 sec science yes 1.8 - 2+" 37002 900 sec sky no +11" in RA 37003 900 sec science yes 2.5 - 3+", S/N < 2, declared bad 37004 1200 sec science yes 2.5 - 3+", S/N < 2, declared bad 05:06Z - Plate 2387 (cartridge 6): Segue. An fk5 was needed to acquire the field. Seeing is still lousy (2.5"). The first exposure (37007) had S/N < 2 for b1 and so it was declared bad. The second exposure (37008) was accidentally ended while trying to increase the exposure time (endStare was clicked rather than resumeStare) and so it was also declared bad. Heavy clouds rolled in shortly after the 3rd exposure was started. The exposure was paused in the hope that the clouds would soon depart but after 30 minutes this exposure was aborted. We sat and waited for the clouds to thin. Around 07:21Z the clouds started to clear, the seeing improved to 2" and we started taking data again. We got 2 good exposures and took a midCal before the clouds became problematic again. The next 2 exposures (37013, 37014) had S/N < 2 for b1 so we declared them bad. The next exposure was aborted due to clouds, rising humidity and high wind and we closed again around 10:00Z. 10:00Z - Enclosure on, doors partially closed, fans still on. The clouds started to thin around 10:45Z but the wind speed was still too high to open. 12:15Z - Wind speed still too high. We give up. Building closed, fans off. endNight rejoiced - sort of. For some reason, endNight did not find any cloud camera images. Viktor ran endNight as himself but we're not sure if this had anything to do with the problem. If anyone cares about the cloud camera images for tonight, they'll have to be copied manually. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 6009 22:40Z 22:56Z 100 O -177.87 -173.67 42 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): YES QA Procedures Done (y/n): YES UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 36993, plate 2144 ------- 22:35 36993 10.0 flat 22:38 36994 2.0 arc 22:41 36995 2.1 arc 22:43 36996 2.1 arc ----- sequence 36997, plate -9999 ------- 23:42 36997 0.0 bias 00:28 36998 0.0 bias ----- sequence 36999, plate 1245 ------- 3.9 2.5 3.6 4.8 not done 03:21 36999 10.0 flat 03:23 37000 2.0 arc 03:44 37001 900.1 target science 04:02 37002 900.1 target sky 04:21 37003 900.1 target science - declared bad - low S/N 04:51 37004 1200.1 target science - declared bad - low S/N ----- sequence 37005, plate 2387 ------- 5.4 12.8 14.4 10.0 not done 05:15 37005 10.0 flat 05:18 37006 2.0 arc 05:41 37007 1200.1 target declared bad - low S/N 05:48 37008 247.1 target declared bad - low S/N 07:43 37009 1200.1 target 08:11 37010 1500.1 target 08:13 37011 10.0 flat 08:16 37012 2.0 arc 08:45 37013 1500.1 target declared bad - low S/N 09:18 37014 1800.2 target declared bad - low S/N ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 03:28Z 8 1245 8.91 -0.0009 56.85 0.0036 187.41 -0.0049 1.000250 05:21Z 6 2387 106.57 0.0028 53.16 0.0037 -99.44 -0.0154 1.000200 07:24Z 6 2387 106.03 0.0063 78.23 0.0032 -81.83 0.0168 1.000150 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 03:29Z 8 1245 1.00025 -2469 -783 100 9 56.9 3.8 14 265 ---- 2.0" 05:22Z 6 2387 1.00020 -1975 -289 175 107 53.2 2.7 8 290 ---- 2.6" 07:24Z 6 2387 1.00015 -1481 34 140 106 78.1 2.5 9 341 ---- 2.0" ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:33Z 39 18 16 315 (NW) 494 - 23:03Z 40 18 12 293 (WNW) 579 - 23:35Z 41 18 16 303 (WNW) 615 - partly cloudy 01:11Z 39 17 13 268 (W) 814 - mostly cloudy 01:41Z 39 17 15 276 (W) 894 - 02:13Z 39 17 9 267 (W) 984 - 02:46Z 39 18 13 268 (W) 1067 - partly cloudy 03:18Z 39 18 14 278 (W) 1143 - clear 03:48Z 38 18 15 268 (W) 1077 - partly cloudy 04:18Z 37 18 10 270 (W) 1064 - clear 04:49Z 36 19 9 272 (W) 1050 - 05:20Z 36 19 17 297 (WNW) 983 - partly cloudy 05:51Z 37 18 15 326 (NW) 1038 - mostly cloudy 06:23Z 38 18 22 331 (NNW) 813 - 06:55Z 37 18 11 327 (NNW) 717 - 07:27Z 36 18 20 328 (NNW) 752 - partly cloudy 07:58Z 35 18 17 327 (NNW) 669 - 08:28Z 34 18 14 339 (NNW) 657 - 08:59Z 32 21 7 355 (NNW) 560 - 09:29Z 30 24 10 355 (NNW) 480 - 10:00Z 30 24 11 15 (NNE) 457 - mostly cloudy 10:31Z 29 22 3 318 (NW) 275 - partly cloudy 11:02Z 28 19 23 7 (N) 240 - 11:33Z 27 14 17 357 (NNW) 334 - 12:03Z 26 14 18 35 (NE) 391 - 12:34Z 26 14 7 17 (NNE) 367 - ---=== TELESCOPE STATUS ===--- 23:20Z Enclosure doors opened, fans and blowers on, louvers opened. 00:45Z Enclosure off. 01:20Z Enclosure on, doors partially closed, fans still on (clouds). 03:00Z Enclosure off. 10:00Z Enclosure on, doors partially closed, fans still on (high wind, clouds, rising humidity). 12:15Z Doors closed, fans and blowers off, louvers closed. Telescope stowed at 30 degrees with cartridge 6 (plate 2387) mounted. Counterweights at 245. Spectro auto-fill connected. Spectro dewar 13:02 196 lb, 22 psi Imager dewar 13:02 201 lb, 20 psi No interlocks bypassed. ---=== SOFTWARE USED ===--- IOP/SOP: v4_12_0 Watcher: v2_31_0 MCP: v5_26_2 TPM: tpm_v3_3_0 AstroDa: v15_10 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_09 SoS: v5_1_0 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.8410 -5.8440 Axial B -5.6830 -5.6770 Axial C 0.9320 0.9420 Trans D -9.9960 -9.9940 Lateral E 10.3759 10.3759 Lateral F 11.9507 11.9380 GALILS Commanded: 4700. -2450. 3000. -7800. 3300. 3000. Actual: 4700. -2450. 3000. -7800. 3300. 3000. SETMIR VALUES PriDesOrient: 0.00 -1.68 7.05 385.25 92.02 PriOrient: 0.00 -1.52 7.02 384.99 93.81 SECONDARY: ---------- Focus: 0.00 Air Temp.: 5.3 Alt.: 30.356434 MIGS TONIGHT NOMINAL Axial A 1.5410 1.5420 Axial B 1.0610 1.0470 Axial C 1.1730 1.1650 Trans D -0.1190 -0.1240 GALILS Commanded: 1593413. 1603624. 1554282. -6800. -5900. Actual: 1593450. 1603650. 1554300. -6800. -5900. SETMIR VALUES SecDesOrient: 1257.00 15.87 0.00 -50.00 136.76 SecOrient: 1256.98 15.87 0.00 -49.55 136.73 ---=== PROBLEMS IN DETAIL ===--- Spectro Focus Details: sp1 sp2 Coll_motor_A 6751 -7207 Coll_motor_B 6753 -10490 Coll_motor_C 6738 -9825 im> caget tpm_TM_PMSSAvg tpm_TM_TelairAvg.VALA tpm_TM_PMSSAvg 2.720 tpm_TM_TelairAvg.VALA 4.575 b1 r1 b2 r2 Mean Offset .10 .09 -.01 -.06 Recommended Piston -860 580